The futures price of gold is $1,000. Futures contracts are for 100 ounces of gold, and the margin requirement is $3,000 a contract. The maintenance market requirement is $1,500. A speculator expects the price of gold to rise and enters into a contract to buy gold.

a. How much must the speculator initially remit?
b. If the futures price of gold rises to $1,005, what is the profit and return on the position?
c. If the futures price of gold declines to $998, what is the loss on the position?
d. If the futures price declines to $984, what must the speculator do?
e. If the futures price continues to decline to $982, how much does the speculator have in the account?
Answer #1

1.
=3000

2.
Profit=100*(1005-1000)=500
Return=500/3000=16.67%

3.
Loss=100*(1000-998)=200

4.
Margin will become=3000+1000*(984-1000)=1400
As this is less than maintenance margin, speculator will have to put additional funds

P.S.: I am not allowed to answer more than 4 questions